Can someone help solve this problem for me?
What is 48.3% of 1,500?
Can you show me how you get your answer?
Thanks,
Sher

Answer by Nate(3500)

48.3% = 48.3/10 = 483/1000 = 0.483
(0.483)(1500)
724.5